given the functions f(x)=x^2+3x+2 and g(x)=x+1, determine a simplified equation for h(x)=f(g(x)).
Answer by ikleyn(52776)

.

h(x) = f(g(x)) = (g(x))^2 + 3*g(x) + 2 = 


     = (x+1)^2 + 3*(x+1) + 2 = (x^2 + 2x + 1) + 3*(x + 1) + 2 = 

                             = x^2 + 2x + 1 + 3x + 3 + 2 = 

                             = now combine like terms = 

                             = x^2 + 5x + 6.    ANSWER

Solved.